TORSIONAL SPECTRUM OF ^{12}CH_{3}^{13}CH_{3}: A 2-STATE FREQUENCY ANALYSIS OF THE TORSIONAL BANDS AND THE ν_{12} VIBRATIONAL FUNDAMENTAL

L. BORVAYEH, \underline{N. MOAZZEN-AHMADI}, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, CANADA; V.-M. HORNEMAN, Department of Physical Sciences, University of Oulu, PO Box 3000, Fin-90014, Oulu, Finland.

The far-infrared spectrum of ^{12}CH_{3}^{13}CH_{3} is studied between 220 − 350 cm^{−1} at an effective resolution of 0.004 − 0.005 cm^{−1} using a Bruker IFS-120 HR . Observation of the torsional fundamental ν_{6} (289 cm^{−1}) and the torsional hot band 2ν_{6} − ν_{6} (255 cm^{−1}), together with the lowest frequency vibrational fundamental ν_{12} (821.5 cm^{−1}), also obtained with the same spectrometer at an effective resolution of 0.0025 cm^{−1}, gives information on the torsional stack of the ground vibrational state and that for v_{12} = 1 state. The frequencies are analysed in terms of a 2-state fit to determine the torsion mediated Coriolis interactions between the torsional stacks. A comparison is made with a 2-state fit of similar data for 12CH_{3}^{12}CH_{3}.